OpenStudy (anonymous):

Solve the equation 2(3^x) = 3(4^x). Give an exact answer and a decimal approximation rounded to one decimal place.

OpenStudy (anonymous):

I think I have to use logs to solve this

OpenStudy (loser66):

first off: moving the alike terms to one side \[2(3^x)=3(4^x)\\\dfrac{1}{3}3^x=\dfrac{1}{2}4^x\] observe: \(\dfrac{1}{3}= 3^{-1}\) and \(\dfrac{1}{2} = 4^{-1/2}\) so that your equation become \[3^{-1}3^x=4^{-1/2}4^x\\3^{x-1}=4^{x-1/2}\] can you handle it from here?

OpenStudy (anonymous):

this is different as to how I was going to handle It, I thought the first step would be to divide both sides by 2 and then solve using logarithms

OpenStudy (loser66):

of course you have to use log to solve, now take ln both sides to get \[ln(3^{x-1})= ln(4^{x-1/2})\\(x-1) ln 3 = (x-1/2)ln4\] so that \[\dfrac{x-1}{x-1/2}=log_3 4\] now solve for x